# revtex4-2Repair the float

### v1

@@ II.1 Single Parameter Learning
To start, let us examine the simpler scenario where there is an error in only one Hubbard model parameter. Specifically, we will look at the case of three quantum dots with an error or deviation only in $\delta\epsilon$. The scenarios involving one and two quantum dots were not considered, as these situations are well understood, and the use of machine learning would be trivial and an overkill. For the scenario where only $\delta\epsilon_{i}\neq 0$, we assume $\delta V_{i,i+1}=\delta t_{i,i+1}=\delta U_{i}=0$. In this model, $\delta\epsilon_{i}\in[-0.5,0.5]$ is chosen from either a Gaussian or a normal distribution with an equal probability. The remaining parameters are kept constant at $V_{i,i+1}=0.2$, $t_{i,i+1}=1$, and $U_{i}=4$. (Unless otherwise stated specifically, we use the nearest neighbor hopping t as the energy unit.)
@@
To further assess the validity of our method, we test whether our predictions for the Hubbard deviations can be used to regenerate the stability diagrams. This is crucial because it is not immediately clear if the collection of charge stability diagrams we feed in can uniquely determine the Hubbard parameters in the first place since there is no ’inverse scattering’ theorem proven for such a scenario. Nonetheless, a match of the charge stability diagrams would indicate how well our predictions capture the physics. Solving for 2 or 3 quantum dots is how this issue is currently addressed experimentally, requiring manual effort and certain restrictions on the parameters. Considerable time is spent periodically to re-tune the dots by determining and then correcting for $\delta\epsilon_{i}$. With our machine learning model, we find we can solve this problem efficiently with high accuracy. Our model, in the case of three quantum dots with only $\delta\epsilon_{i}$ errors, achieves an $R^{2}=0.99996$ with an RMS error of $R_{MS}(\delta\epsilon)=0.00183$ at each site. We are not just finding $\delta\epsilon_{i}$ for individual sites but for the entire position-based function. The results, shown in Fig. 3, illustrate that deviations in gate voltages primarily result in shifts in the phase boundaries, making $\delta\epsilon$ seemingly the easiest parameter to determine. This suggests that, at least for site-specific $\epsilon$ errors within our parameter range, the convolutional neural network can uniquely identify the $\delta\epsilon$ disorder by using merely a series of nearest neighbor charge stability diagrams. This outcome is significant (and utterly non-obvious) as it could enable the automation of re-tuning this single parameter, removing the need for manual intervention, which is not significantly restricted by the smallness of $t$ as mentioned in [10].
@@
However, solving $\delta\epsilon_{i}$ for 3 quantum dots beyond the t range, while beneficial for automation and possibly the speed of determination, does not address a problem that is not already regularly solved manually in the experimental setting. The real machine learning advantage lies in its ability to solve for other unknown parameters and the fact that it can do so with larger numbers of quantum dots at once. Currently, quantum dots are experimentally tuned two or three at a time, with this aggregate tuning deemed sufficient for the whole device [10]. Our method, however, allows many more quantum dots to be tuned at once, providing more accurate and much faster automated tuning, and enabling more parameters to be adjusted. In the case of four quantum dots with errors only in $\delta\epsilon_{i}$ using the same parameter setup, we achieved an $R^{2}=0.99995$ and an average prediction error of $R_{MS}(\delta\epsilon)=0.0029$. This is an impressive result that, unlike previous work, allows four or more quantum dots to be tuned simultaneously. A latter section will demonstrate that this approach is effective for five quantum dots as well, with minimal impact on fidelity. Yet, the capability to tune more dots simultaneously is not the only significant advantage of our method; it also enables tuning of parameters beyond $\delta\epsilon$ that are more challenging (and typically not done manually in the laboratory).
@@
Next, we examine the case of determining site-dependent values of $\delta V_{ij}$, $\delta t_{ij}$, and $\delta U_{i}$, where we assume errors in only one parameter at a time and presume accurate estimations can be achieved for the other parameters. This approach is taken to determine whether our CNN scheme can predict all these parameters or just $\epsilon_{i}$ errors. In a single parameter error scenario with four quantum dots and approximately 40,000 training realizations, we find that $\delta V_{ij}$ can be predicted very accurately with an $R^{2}=0.99995$ and $R_{MS}(\delta V_{ij})=0.0009$ and is shown in Fig. 15. $\tilde{V}_{ij}$ is assumed to be within a range of $[0,0.4]$, for which we use both Gaussian and uniform distributions at random. This demonstrates the CNN’s ability to accurately predict this parameter for each bond uniquely. In the case of single parameter error $\delta t_{ij}$ at each bond, the CNN predicts the deviation with an $R^{2}=0.99995$ and $R_{MS}(\delta t_{ij})=0.0243$, and is shown in Fig. 16. $\tilde{t}_{ij}$ is assumed to be within a range of $[0.1,10]$, selected using a uniform distribution. For single error $\delta U_{i}$, the CNN predicts the deviation with an $R^{2}=0.9988$ and $R_{MS}(\delta U_{i})=0.101$, and is shown in Fig. 17. $\tilde{U}_{i}$ is assumed to be approximately within $[0,8]$, where we alternate between Gaussian or uniform distribution at random, ensuring that the correct sign is maintained. In all cases, and thus for all our Hubbard parameters, the neural network accurately predicts the deviation errors from just a series of charge stability diagrams. This is an important (and rather surprising as well as useful) finding since charge stability diagrams are routinely measured experimentally in quantum dot qubit devices.
@@
Figure 3: 3 quantum dot charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of only $\delta{\epsilon_{i}}$. The root mean squared error in the Hubbard model $\epsilon_{i}$ parameter was $R_{MS}({\delta\epsilon})=0.00183$ with an $R^{2}=0.99998$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. The 2D type CNN was used for this model. The columns of these two subplots correspond to the same representative samples. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the disorder free values respectively), this is similar to the stability diagrams fed into the neural network.
@@
Figure 5: 4 quantum dot charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of only $\delta{\epsilon_{i}}$. The root mean squared error in the Hubbard model parameter was $R_{MS}({\delta\epsilon})=0.0029$ with an $R^{2}=0.99995$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. The 2D type CNN was used for this model. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively), this is similar to the stability diagrams fed into the neural network except the neural network only receives $\mu_{i}=-\mu_{j}$ between nearest neighbors $i=j+1$.
@@ II.2 Many Parameter Learning
In the last section, it was shown that high-fidelity predictions of individual Hubbard parameters can be achieved using just a few stability diagrams, assuming that deviations for all but one Hubbard parameter are small. This approach allows for accurate predictions of the deviation in the remaining parameter. However, in practical experiments, all parameters typically exhibit errors that need to be addressed simultaneously. To meet this challenge, we introduce deviations from the expected values for all our Hubbard parameters. Our machine learning method, leveraging nearest neighbor stability diagrams, is capable of determining all the Hubbard parameters at once, without any initial knowledge of any specific parameter (or no knowledge except a reasonable range). In our investigations involving multiple parameters, we train our neural network with 10,000 to 20,000 training realizations. This is a reduction from the 40,000 realizations used in single-parameter scenarios, motivated by our findings that so many realizations are not necessary for achieving high fidelity and by the increased numerical demand of the simulations when all parameters exhibit disorder. Nonetheless, it is apparent that including more training realizations could further enhance the (already pretty high) fidelity.
@@
To begin, consider the 3 quantum dot case, with a more limited parameter range of $\tilde{\epsilon}_{i}\in[-0.5,0.5]$, $\tilde{V}_{ij}\in[0,0.4]$, $\tilde{t}_{ij}\in[0.01,0.25]$, and $\tilde{U}_{i}\in[3,5]$. To set a scale for the system, $t_{12}=0.125$ is fixed. We find that the system can predict the Hubbard deviations with an $R^{2}=0.9987$ and RMS errors of $R_{MS}({\delta\epsilon_{i}})=0.0128$, $R_{MS}(\delta V_{i,j})=0.0190$, $R_{MS}({\delta t_{i,j}})=0.0119$, and $R_{MS}(\delta U_{i})=0.0358$. Although this prediction is not as precise as in the single parameter regime, it remains very accurate. This outcome strongly suggests that the CNN, when fed with charge stability diagrams, can predict all the Hubbard model parameters simultaneously. The figure for this can be viewed in Fig. 6. An assessment of the actual error within the stability diagrams due to inaccuracies in the predictions was also performed. This assessment is crucial because our plots only display the most probable state, and it is possible that the differences between the predicted and input stability diagrams are not as significant as they appear, potentially being an artifact of this selection process. To explore this, we plot the magnitude of the error in the occupation expectation vectors between the input and predicted Hubbard model parameters as $||\langle\vec{n}_{input}\rangle-\langle\vec{n}_{predicted}\rangle||$ in Fig. 9. It was found that the errors are concentrated at the boundaries, which could be due to the low resolution of the stability diagrams inputted into the neural network. This may also result from the majority of information about the Hubbard parameters in the charge stability diagram being concentrated within the phase boundaries. This suggests that our grid scheme of generating the charge stability diagrams is sub-optimal, and utilizing a more sophisticated edge-finding method to extrapolate the diagrams to a higher resolution could lead to significant improvements. The specific impact of increasing the resolution will be discussed later. Nonetheless, these errors are relatively minor, with the stability diagrams being well reproduced from the predicted Hubbard model parameters.
@@
Transitioning to the main many parameter regime, we expand $\tilde{t}_{ij}\in[0.1,2]$ and establish a scale for the system by fixing $t_{12}=1$. It is crucial to recognize that this main many parameter regime is still more constrained than the single-parameter predictions for $\delta U_{i}$ and $\delta t_{ij}$. Despite a decrease in fidelity within this expanded range, the system still achieves an $R^{2}=0.9977$ with errors of $R_{MS}({\delta\epsilon_{i}})=0.0239$, $R_{MS}(\delta V_{i,j})=0.0164$, $R_{MS}({\delta t_{i,j}})=0.0121$, and $R_{MS}(\delta U_{i})=0.0373$. This shows that increasing our t range, which significantly increases the complexity of the problem, is still able to be handled within some sufficiently small (but still relative to experiment large) range.
@@
Expanding to a larger number of quantum dots, we first examine the case involving 4 coupled quantum dots with deviations in all parameters. We achieved an $R^{2}=0.9961$ with errors of $R_{MS}({\delta\epsilon_{i}})=0.0280$, $R_{MS}({\delta V_{i,j}})=0.0205$, $R_{MS}({\delta t_{i,j}})=0.0193$, and $R_{MS}({\delta U_{i}})=0.0512$. This represents only a slight reduction in the fidelity of our predictions, indicating that the tuning approach remains effective for a larger number of quantum dots. The charge stability diagrams are illustrated in Fig. 10, with the corresponding stability diagram errors depicted in Fig. 12.
@@
Our method also performs well in the case of 5 quantum dots with deviations across all parameters. We managed to achieve an $R^{2}=0.995$ with errors of $R_{MS}({\delta\epsilon_{i}})=0.0321$, $R_{MS}(\delta V_{i,j})=0.0231$, $R_{MS}({\delta t_{i,j}})=0.0254$, and $R_{MS}(\delta U_{i})=0.0574$. Once again, this shows only a minor decrease in prediction fidelity, underscoring that the tuning method is viable for an increasing number of quantum dots. The charge stability diagrams for this scenario are presented in Fig. 11, with the respective errors shown in Fig. 13. These results suggest that the machine learning scheme can automatically tune as many quantum dots as one has computational resources for training data generation, and it can be accomplished with a relatively simple neural network. Given that the Hubbard model is a strongly interacting system, the generation of training data is the primary bottleneck by a wide margin, while machine learning resources, by comparison, do not significantly scale with the number of dots. Ideally, the number of quantum dots one needs to tune simultaneously depends on the effect neighboring quantum dots have on the local state of the system. Bearing this in mind, it is not seemingly necessary when tuning a large system to arbitrarily scale up the number of dots tuned simultaneously. In fact, the 5-dot system achieved in our work should describe well all current and near-future quantum dot qubit platforms. When used on actual experiments, the effective Hubbard model parameters will have to be generated first by initially fitting to the experimental data. The machine learning procedure outlined in our work then can be effectively applied to the experiments.
@@
Our neural network faces its greatest challenge in accurately predicting the Hubbard model parameters over a very broad $\tilde{t}_{ij}$ range. The curvature induced by $t$ in the charge stability diagrams might complicate the network’s task of discerning other parameters, likely necessitating more training data due to the system’s increased complexity. When $\tilde{t}$ ranges from $\tilde{t}\in[0.1,10]$, significant errors are found in the predictions of parameters other than $\tilde{t}$. However, this issue remains solvable, as even when predictions for other parameters "fail," the disordered $\tilde{t}$ can still be predicted with high accuracy. For scenarios requiring a large $t$ range, we recommend a sequential procedure, initially generating a prediction for $\tilde{t_{i}}$ using the full range, then creating training data and training a new neural network within a narrower $t$ range centered around the initial prediction. This approach reduces the dominance of $t$ by focusing the training data within a probable range. Alternatively, generating a significantly larger volume of training data from the outset is an option, though it may be less efficient. Additionally, employing a standard scaling operation, such as those available in the scikit-learn package [17], proves useful when dealing with output parameters of substantially different scales. This technique ensures that the neural network does not prioritize only predicting $\tilde{t}$ by mapping all parameters in an invertible manner to a Gaussian distribution with a standard deviation of 1 and a mean of 0, facilitating a more balanced and accurate prediction process across all parameters. We note that the hopping amplitude t represents the quantum fluctuation (and consequently also the 2-qubit exchange coupling $J=t^{2}/U$), and it is interesting that the machine learning procedure becomes more demanding with increasing quantum fluctuations in the strong-coupling quantum dot qubit problem.
@@
To demonstrate the effects of resolution, we apply our method to a 3 quantum dot system within the restricted $t$ range ($\tilde{t}_{ij}\in[0.01,0.25]$), but with double the resolution—and therefore the number of steps—in the $\mu$ and $\mu_{i}^{j}$ parameters. This approach yielded an $R^{2}=0.9984$ with errors of $R_{MS}(\delta\epsilon_{i})=0.0109$, $R_{MS}(\delta V_{i,j})=0.0160$, $R_{MS}(\delta t_{i,j})=0.0089$, and $R_{MS}(\delta U_{i})=0.0423$. These results mark an improvement in the RMS errors compared to the case with lower resolution. Additionally, employing standard scaling further enhances performance, leading to an $R^{2}=0.9972$ and $R_{MS}(\delta\epsilon_{i})=0.0110$, $R_{MS}(\delta V_{i,j})=0.0113$, $R_{MS}(\delta t_{i,j})=0.0045$, and $R_{MS}(\delta U_{i})=0.0347$. Notably, the error $R_{MS}({\delta t_{i,j}})$ is about half that of the lower resolution case without standard scaling. It is is also worth noting that since $R^{2}$ is affected by standard scaling, it is not directly comparable with the previous results.
@@
Figure 7: 3 quantum dot charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of disorder in all parameters. The root mean squared error in the Hubbard model parameters was $R_{MS}({\delta\epsilon_{i}})=0.0239$, $R_{MS}({\delta V_{i,j}})=0.0164$, $R_{MS}({\delta t_{i,j}})=0.0121$, $R_{MS}({\delta U_{i}})=0.0373$ with an $R^{2}=0.9977$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the their disorder free values respectively), this is the data input into the machine learning model.

### v2

@@ II.1 Single Parameter Learning
To start, let us examine the simpler scenario where there is an error in only one Hubbard model parameter. Specifically, we will look at the case of three quantum dots with an error or deviation only in $\delta\epsilon$. The scenarios involving one and two quantum dots were not considered, as these situations are well understood, and the use of machine learning would be trivial and an overkill. For the scenario where only $\delta\epsilon_{i}\neq 0$, we assume $\delta V_{i,i+1}=\delta t_{i,i+1}=\delta U_{i}=0$. In this model, $\delta\epsilon_{i}\in[-0.5,0.5]$ is chosen from either a Gaussian or a normal distribution with an equal probabilities to increase the complexity in the data and demonstrate our methods robustness). When applicable all deviations will use both types of distributions except $\delta t_{i,i+1}$ which will only use uniform. The remaining parameters are kept constant at $V_{i,i+1}=0.2$, $t_{i,i+1}=1$, and $U_{i}=4$. (Unless otherwise stated specifically, we use the nearest neighbor hopping t as the energy unit.)
@@
To further illustrate the validity of our method, we test whether our predictions for the Hubbard deviations can be used to regenerate the stability diagrams. This is crucial because it is not immediately clear if the collection of charge stability diagrams we feed in can uniquely determine the Hubbard parameters in the first place since there is no ’inverse scattering’ theorem proven for such a scenario. Nonetheless, a match of the charge stability diagrams would illustrate how well our predictions capture the physics. Solving for 2 or 3 quantum dots is how this issue is currently addressed experimentally, requiring manual effort and certain restrictions on the parameters. Considerable time is spent periodically to re-tune the dots by determining and then correcting for $\delta\epsilon_{i}$. With our machine learning model, we find we can solve this problem efficiently with high accuracy. Our model, in the case of three quantum dots with only $\delta\epsilon_{i}$ errors, achieves an $R^{2}=0.99996$ with an RMS error of $R_{MS}(\delta\epsilon)=0.00183$ at each site. Both $R^{2}$ and $R_{MS}$ metrics presented are in terms of the Hubbard parameter deviations to assess the validity of our method. We are not just finding $\delta\epsilon_{i}$ for individual sites but for the entire position-based function. The results, shown in Fig. 4, illustrate that deviations in gate voltages primarily result in shifts in the phase boundaries, making $\delta\epsilon$ seemingly the easiest parameter to determine. This suggests that, at least for site-specific $\epsilon$ errors within our parameter range, the convolutional neural network can uniquely identify the $\delta\epsilon$ disorder by using merely a series of nearest neighbor charge stability diagrams. This outcome is significant as it could enable the automation of re-tuning this single parameter, removing the need for manual intervention, which is not significantly restricted by the smallness of $t$ as mentioned in [14].
@@
However, solving $\delta\epsilon_{i}$ for 3 quantum dots beyond the t range, while beneficial for automation and possibly the speed of determination, does not address a problem that is not already regularly solved manually in the experimental setting. The real machine learning advantage lies in its ability to solve for other unknown parameters and the fact that it can do so with larger numbers of quantum dots at once. Currently, quantum dots are experimentally tuned two or three at a time, with this aggregate tuning deemed sufficient for the whole device [14]. Our method, however, allows many more quantum dots to be tuned at once, providing more accurate and much faster automated tuning, and enabling more parameters to be adjusted. In the case of four quantum dots with errors only in $\delta\epsilon_{i}$ using the same parameter setup (see Fig. A4), we achieved an $R^{2}=0.99995$ and an average prediction error of $R_{MS}(\delta\epsilon)=0.0029$. This is an impressive result that, unlike previous work, allows four or more quantum dots to be tuned simultaneously. A latter section will demonstrate that this approach is effective for five quantum dots as well, with minimal impact on fidelity. Yet, the capability to tune more dots simultaneously is not the only significant advantage of our method; it also enables tuning of parameters beyond $\delta\epsilon$ that are more challenging (and typically not done manually in the laboratory).
@@
Next, we examine the case of determining site-dependent values of $\delta V_{ij}$, $\delta t_{ij}$, and $\delta U_{i}$, where we assume errors in only one parameter at a time and presume accurate estimations can be achieved for the other parameters. This approach is taken to determine whether our CNN scheme can predict all these parameters or just $\epsilon_{i}$ errors. In a single parameter error scenario with four quantum dots and approximately 40,000 training realizations, we find that $\delta V_{ij}$ can be predicted very accurately with an $R^{2}=0.99995$ and $R_{MS}(\delta V_{ij})=0.0009$ and is shown in Fig. A5. $\tilde{V}_{ij}$ is assumed to be within a range of $[0,0.4]$. This demonstrates the CNN’s ability to accurately predict this parameter for each bond uniquely. In the case of single parameter error $\delta t_{ij}$ at each bond, the CNN predicts the deviation with an $R^{2}=0.99995$ and $R_{MS}(\delta t_{ij})=0.0243$, and is shown in Fig. A6. $\tilde{t}_{ij}$ is assumed to be within a range of $[0.1,10]$. For single error $\delta U_{i}$, the CNN predicts the deviation with an $R^{2}=0.9988$ and $R_{MS}(\delta U_{i})=0.101$, and is shown in Fig. A7. $\tilde{U}_{i}$ is assumed to be approximately within $[0,8]$, ensuring that the correct sign is maintained. In all cases, and thus for all our Hubbard parameters, the neural network accurately predicts the deviation errors from just a series of charge stability diagrams. This is an important finding since charge stability diagrams are routinely measured experimentally in quantum dot qubit devices.
@@
Figure 4: 3 quantum dot charge stability diagrams for input and expected measurement outcomes from Hubbard model parameters for the prediction of only $\delta{\epsilon_{i}}$. The root mean squared error in the Hubbard model $\epsilon_{i}$ parameter was $R_{MS}({\delta\epsilon})=0.00183$ with an $R^{2}=0.99998$. In both plots the first row is the input charge stability diagrams, namely the most probable state for a chemical potential configuration, and the second row is the expected charge stability diagram given the prediction of the Hubbard model parameters. The error-free model parameters were set to $U=4$, $t=1$ and $V_{i,i+1}=0.2$. The 2D type CNN was used for this model. The columns of these two subplots correspond to the same randomly selected representative validation samples. (a) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where $\mu_{1}$ and $\mu_{3}$ are independently varied. (b) Most probable state for input Hubbard parameters (1st row) and predicted Hubbard parameters (2nd row) where the chemical potential at each site is $\vec{\mu}=[\mu_{1},\cdots,\mu_{n}]$ with our plot having axis $\mu_{1}=-\mu_{3}$ (rescaled by $\langle U\rangle$ and shifted by $\mu$) vs. $\mu_{k\neq 1,3}=\mu$ (rescaled by $W=\frac{1}{L}\left(\sum_{i}\langle U_{i}\rangle+\langle V_{i,i+1}\rangle\right)$ where $\langle U\rangle=4$ and $\langle V_{i,i+1}\rangle=0.2$ are the disorder free values respectively), this is similar to the stability diagrams fed into the neural network.
@@ II.2 Many Parameter Learning
Building on high-fidelity predictions of the previous section, in practical experiments, all parameters typically exhibit errors that need to be addressed simultaneously. To meet this challenge, we introduce deviations from the expected values for all our Hubbard parameters. Our machine learning method, leveraging nearest neighbor stability diagrams, is capable of determining all the Hubbard parameters at once, without any initial knowledge of any specific parameter beyond a very rough range/scale necessary for generating training data. In our investigations involving multiple parameters, we train our neural network with 10,000 to 20,000 training realizations. This is a reduction from the 40,000 realizations used in single-parameter scenarios, motivated by our findings that so many realizations are not necessary for achieving high fidelity and by the increased numerical demand of the simulations when all parameters exhibit disorder. Nonetheless, it is apparent that including more training realizations could further enhance the (already pretty high) fidelity.
@@
To begin, consider the 3 quantum dot case, with a more limited parameter range of $\tilde{\epsilon}_{i}\in[-0.5,0.5]$, $\tilde{V}_{ij}\in[0,0.4]$, $\tilde{t}_{ij}\in[0.01,0.25]$, and $\tilde{U}_{i}\in[3,5]$. To set a scale for the system, $t_{12}=0.125$ is fixed. We find that the system can predict the Hubbard deviations with an $R^{2}=0.9987$ and RMS errors of $R_{MS}({\delta\epsilon_{i}})=0.0128$, $R_{MS}(\delta V_{i,j})=0.0190$, $R_{MS}({\delta t_{i,j}})=0.0119$, and $R_{MS}(\delta U_{i})=0.0358$. Although this prediction is not as precise as in the single parameter regime, it remains very accurate. The figure for this can be viewed in Fig. 6. An assessment of the actual error within the stability diagrams due to inaccuracies in the predictions was also performed. This assessment is crucial because our plots only display the most probable state, and it is possible that the differences between the predicted and input stability diagrams are not as significant as they appear, potentially being an artifact of this selection process. To explore this, we plot the magnitude of the error in the occupation expectation vectors between the input and predicted Hubbard model parameters as $||\langle\vec{n}_{input}\rangle-\langle\vec{n}_{predicted}\rangle||$ in Fig. 8 and Fig. 5. It was found that the errors are concentrated at the boundaries, which could be due to the low resolution of the stability diagrams inputted into the neural network. This may also result from the majority of information about the Hubbard parameters in the charge stability diagram being concentrated within the phase boundaries. This suggests that our grid scheme of generating the charge stability diagrams is sub-optimal, and utilizing a more sophisticated edge-finding method to extrapolate the diagrams to a higher resolution could lead to significant improvements. The specific impact of increasing the resolution will be discussed later. Nonetheless, these errors are relatively minor, with the stability diagrams being well reproduced from the predicted Hubbard model parameters.
@@
Transitioning to the main many parameter regime, we expand $\tilde{t}_{ij}\in[0.1,2]$ and establish a scale for the system by fixing $t_{12}=1$. It is crucial to recognize that this main many parameter regime is still more constrained than the single-parameter predictions for $\delta U_{i}$ and $\delta t_{ij}$. Despite a decrease in fidelity within this expanded range, the system still achieves an $R^{2}=0.9977$ with errors of $R_{MS}({\delta\epsilon_{i}})=0.0239$, $R_{MS}(\delta V_{i,j})=0.0164$, $R_{MS}({\delta t_{i,j}})=0.0121$, and $R_{MS}(\delta U_{i})=0.0373$ (see Fig. A1 and Fig. A2). This shows that increasing our t range, which significantly increases the complexity of the problem, is still able to be handled within some sufficiently small (but still relative to experiment large) range.
@@
Expanding to a larger number of quantum dots, we first examine the case involving 4 coupled quantum dots with deviations in all parameters. We achieved an $R^{2}=0.9961$ with errors of $R_{MS}({\delta\epsilon_{i}})=0.0280$, $R_{MS}({\delta V_{i,j}})=0.0205$, $R_{MS}({\delta t_{i,j}})=0.0193$, and $R_{MS}({\delta U_{i}})=0.0512$. This represents only a slight reduction in the fidelity of our predictions, indicating that the tuning approach remains effective for a larger number of quantum dots. The charge stability diagrams are illustrated in Fig. A8, with the corresponding stability diagram errors depicted in Fig. A3.
@@
Our method also performs well in the case of 5 quantum dots with deviations across all parameters. We managed to achieve an $R^{2}=0.995$ with errors of $R_{MS}({\delta\epsilon_{i}})=0.0321$, $R_{MS}(\delta V_{i,j})=0.0231$, $R_{MS}({\delta t_{i,j}})=0.0254$, and $R_{MS}(\delta U_{i})=0.0574$. Once again, this shows only a minor decrease in prediction fidelity, underscoring that the tuning method is viable for an increasing number of quantum dots. The charge stability diagrams for this scenario are presented in Fig. 7, with the respective errors shown in Fig. 9. These results suggest that the machine learning scheme can automatically tune as many quantum dots as one has computational resources for training data generation, and it can be accomplished with a relatively simple neural network. Given that the Hubbard model is a strongly interacting system, the generation of training data is the primary bottleneck by a wide margin, while machine learning resources, by comparison, do not significantly scale with the number of dots. Ideally, the number of quantum dots one needs to tune simultaneously depends on the effect neighboring quantum dots have on the local state of the system. Bearing this in mind, it is not seemingly necessary when tuning a large system to arbitrarily scale up the number of dots tuned simultaneously see Section II.4.
@@
Our neural network faces its greatest challenge in accurately predicting the Hubbard model parameters over a very broad $\tilde{t}_{ij}$ range. The curvature induced by $t$ in the charge stability diagrams might complicate the network’s task of discerning other parameters, likely necessitating more training data due to the system’s increased complexity. When $\tilde{t}$ ranges from $\tilde{t}\in[0.1,10]$, significant errors are found in the predictions of parameters other than $\tilde{t}$. However, this issue remains solvable, as even when predictions for other parameters "fail," the disordered $\tilde{t}$ can still be predicted with high accuracy. For scenarios requiring a large $t$ range, we recommend a sequential procedure, initially generating a prediction for $\tilde{t_{i}}$ using the full range, then creating training data and training a new neural network within a narrower $t$ range centered around the initial prediction. This approach reduces the dominance of $t$ by focusing the training data within a probable range. Alternatively, generating a significantly larger volume of training data from the outset is an option, though it may be less efficient. Additionally, employing a standard scaling operation, such as those available in the scikit-learn package [37], proves useful when dealing with output parameters of substantially different scales. This technique ensures that the neural network does not prioritize only predicting $\tilde{t}$ by mapping all parameters in an invertible manner to a Gaussian distribution with a standard deviation of 1 and a mean of 0, facilitating a more balanced and accurate prediction process across all parameters. We note that the hopping amplitude t represents the quantum fluctuation (and consequently also the 2-qubit exchange coupling $J=4t^{2}/U$), and it is interesting that the machine learning procedure becomes more demanding with increasing quantum fluctuations in the strong-coupling quantum dot qubit problem.
@@
To demonstrate the effects of resolution, we apply our method to a 3 quantum dot system within the restricted $t$ range ($\tilde{t}_{ij}\in[0.01,0.25]$), but with double the resolution—and therefore the number of steps—in the $\mu$ and $\mu_{i}^{j}$ parameters. This approach yielded an $R^{2}=0.9984$ with errors of $R_{MS}(\delta\epsilon_{i})=0.0109$, $R_{MS}(\delta V_{i,j})=0.0160$, $R_{MS}(\delta t_{i,j})=0.0089$, and $R_{MS}(\delta U_{i})=0.0423$. These results mark an improvement in the RMS errors compared to the case with lower resolution. Additionally, employing standard scaling further enhances performance, leading to an $R^{2}=0.9972$ and $R_{MS}(\delta\epsilon_{i})=0.0110$, $R_{MS}(\delta V_{i,j})=0.0113$, $R_{MS}(\delta t_{i,j})=0.0045$, and $R_{MS}(\delta U_{i})=0.0347$ (see Fig. A9). Notably, the error $R_{MS}({\delta t_{i,j}})$ is about half that of the lower resolution case without standard scaling. It is is also worth noting that since $R^{2}$ is affected by standard scaling, it is not directly comparable with the previous results.